OpenStudy (anonymous):

how to find the area between x-axis and g(x) from x=0 to x=pi g(x) = xcos(x/2)

OpenStudy (dumbcow):

\[A = \int\limits_{0}^{\pi}x \cos(x/2) dx\] u = x/2 du = 1/2 dx \[=4\int\limits_{0}^{\pi}u \cos(u) du\] then use integration by parts u = u dv = cos u du = du v = sin u